OpenStudy (baldymcgee6):

For each of the following functions determine (f(x+h)-f(x)) / h and express your answer in simple form. a) f(x)= sqrt(3+4x)

OpenStudy (baldymcgee6):

@ganeshie8

OpenStudy (baldymcgee6):

first principles? ish?

ganeshie8 (ganeshie8):

yea but its not exactly, theres no limit thingy :P put the expression and try to rationalize numerator. i see h would get cancel out may be

OpenStudy (baldymcgee6):

I don't know how to setup the expression

ganeshie8 (ganeshie8):

okay! f(x) is already given. to setup the expression, we need f(x+h) right ?

ganeshie8 (ganeshie8):

\(f(x) = \sqrt{3+4x}\) \(f(x+h) = \sqrt{3+4(x+h)}\)

OpenStudy (baldymcgee6):

okay right, thank you.. i should get it from here

ganeshie8 (ganeshie8):

i just replaced x with (x+h)

ganeshie8 (ganeshie8):

oh great ! put the expression and rationalize numerator. h in the bottom cancels... . good luck :)

OpenStudy (anonymous):

After rationalizing you should obtain \[ \frac{4}{\sqrt{4 (h+x)+3}+\sqrt{4 x+3}} \]

OpenStudy (anonymous):

If h goes to zero, you get \[ \frac{2}{\sqrt{4 x+3}} \] Which is the derivative f(x)